A manufacturing company is examining a proposed project to increase productivity. This project involves buying a machine that will have a cost of $32561 and will have a useful life of 7 years. At the end of the machine's useful life, the company expects to sell it for $2,500. The annual benefits are going to be $22,000 for the first year but will increase by $700 each year until the machine is no longer useful. The estimated operating and maintenance costs are going to be $12,600 for the first year but will increase by $1,000 each year after that. The analysis will use an interest rate of 8%. Calculate this project's net present value or net present worth.
1
Expert's answer
2020-02-27T09:41:40-0500

"NPV = -32,561 + \frac{22,000 - 12,600}{1 + 0.08} + \frac{22,700 - 13,600}{1.08^2} + \frac{23,400 - 14,600}{1.08^3} + \frac{24,100 - 15,600}{1.08^4} + \frac{24,800 - 16,600}{1.08^5} + \frac{25,500 - 17,600}{1.08^6} + \frac{26,200 + 2,500 - 18,600}{1.08^7} = 13,630.34."